# Efficient Posterior Sampling For Diverse Super-Resolution with Hierarchical VAE Prior

### 14 Dec 2023 v3

@@ 1 INTRODUCTION
Image super-resolution is the task of generating a high-resolution (HR) image $\bm{x}$ corresponding to a low-resolution (LR) observation $\bm{y}$. A typical approach for image super-resolution is to train a deep neural network in a supervised fashion to map a LR image to its HR counterpart (see [Lepcha et al., 2022] for an extensive review). Despite impressive performances, those regression based methods are fundamentally limited by their lack of diversity. Indeed, there might exist many plausible HR solutions associated with one LR observation, but regression based methods only provide one of those solutions.
@@
Starting from the seminal work of  [Lugmayr et al., 2020], many approaches proposed to train conditional generative models such as conditional normalizing flow or conditional variational autoencoders in order to model the posterior distribution of the super-resolution problem. We refer to those methods as direct methods, as they only require one network function evaluation (NFE) to generate one sample.
With the recent development of score-based generative models (also known as denoising diffusion models) [Ho et al., 2020, Song et al., 2021], posterior sampling methods based on conditional denoising diffusion models are now able to produce high-quality samples outperforming previous direct methods [Choi et al., 2021, Chung et al., 2022, Kawar et al., 2022]. However, denoising diffusion methods are limited by their computationally expensive sampling process, as they require numerous network function evaluations to generate one super-resolved sample. In the following, we classify those methods as iterative methods.
@@
Our approach is based on reusing a pretrained hierarchical variational autoencoder (HVAE) [Sønderby et al., 2016, Kingma et al., 2016]. HVAE models are able to generate high-quality images by relying on an expressive sequential generative model [Vahdat and Kautz, 2020, Child, 2021, Hazami et al., 2022]. By associating one latent variable subgroup to each residual block of a generative network, HVAE models are able to learn compact high-level representations of the data, and they can generate new samples efficiently, with only one evaluation of the generative network.
The fast sampling time and the expressivity of HVAE models make them suitable candidates for efficient posterior sampling. In this work, we exploit a pretrained VD-VAE model [Child, 2021]. In order to repurpose VD-VAE generative model for image super-resolution, we train a low-resolution encoder to encode LR images in the latent space of the VD-VAE model. By combining the LR encoder with the VD-VAE generative model, we can produce a sample with only one (autoencoder) network evaluation. By adopting a stochastic model for the LR encoder, our method can generate diverse samples from the posterior distribution, as illustrated in Figure 1. We show that the LR encoder can be trained with reasonable computational resources by exploiting the VD-VAE original (HR) encoder to generate labels for training the LR-encoder, and by sharing weights between the LR encoder and VD-VAE generative model. We evaluate our method on super-resolution of face images, with upscaling factor ×$\times 4$ and ×$\times 8$ and demonstrate that it reaches sample quality on par with sequential methods, while being significantly faster (>×500$\times$ 500> × 500).
@@ 2 HIERARCHICAL VAE
A VAE also provides an inference model (encoder) $q_{\phi}\left(\bm{z}|\bm{x}\right)$, trained to match the intractable model posterior $p_{\theta}\left(\bm{z}|\bm{x}\right)$ [Kingma and Welling, 2013]. In order to define expressive models, both the generative model $p_{\theta}\left(\bm{z},\bm{x}\right)$ and the encoder $q_{\phi}\left(\bm{z}|\bm{x}\right)$ are parameterized by neural networks, whose weights are respectively parameterized with $\theta$ and $\phi$.
@@ 3 SUPER-RESOLUTION WITH HVAE
Super-resolution model It has been observed in several works that the low-frequency information of images generated by HVAE model where mostly controlled by the low-resolution latent variable, at the beginning of the hierarchy [Vahdat and Kautz, 2020, Child, 2021, Havtorn et al., 2021]. Preliminary experiments, detailed in appendix A validate those previous observations. Hence, for a large enough number of latent groups $k$, samples from $p_{\theta}\left(\bm{x}|\bm{x}_{<k}\right)$ share the same low-frequency information. As a consequence, all the samples from $p_{\theta}\left(\bm{x}|\bm{x}_{<k}\right)$ are consistent to a LR image $\bm{y}$ (up to a small error). This motivates us to define the following super-resolution model:
@@
The criterion (13) was introduced by [Harvey et al., 2022], who demonstrated that minimizing (13) is equivalent to maximizing a lower-bound of the super-resolution conditional log-likelihood on the training dataset, and that under additional assumptions on the pretrained HVAE model, one can reach optimal performance by only training the low-resolution encoder $q_{\psi}\left(\bm{z}_{<k}|\bm{y}\right)$. In practice, the KL divergence within the training criterion (13) can be decomposed into a sum of KL divergence on each latent subgroup:
@@ 4.1 Experimental settings
We test our super-resolution method CVDVAE on the FFHQ dataset [Karras et al., 2019], with images of resolution $256\times 256$. We experiment on 2 upscaling factors: ×$\times 4$ ($64\times 64\rightarrow 256\times 256$) and ×$\times 8$ ($32\times 32\rightarrow 256\times 256$). The low resolution images are initially downscaled by applying an antialiasing kernel followed by a bicubic interpolation.
Compared methods We compare CVDVAE with a conditional normalizing flow (HCFlow) [Liang et al., 2021], a conditional diffusion model (SR3) [Saharia et al., 2021], and a method that add guidance to a non-conditional diffusion model at inference (DPS) [Chung et al., 2022]. We retrain HCFlow on FFHQ256 using the official implementation. For DPS, we also reuse the official implementation with the available pretrained model, which was trained on FFHQ. For SR3, since no official implementation is available, we used an open-source (non-official) implementation [Jiang, 2022], and we trained a model on FFHQ. When training SR3, we found that a color shift [Deck and Bischoff, 2023] was responsible for important reconstruction errors. To compensate this weakness of the method, we project the super-resolved image on the space of consistent solutions at inference as proposed in [Bahat and Michaeli, 2020]. For fair comparison, we retrained both HCFlow and SR3 with the same computational budget as for our low-resolution encoder. For HCFlow and CVDVAE, we set the temperature of the latent variables at $\tau=0.8$ during sampling.
@@
Evaluation metrics The general quality of the super-resolved images is evaluated using the blind Image quality metric BRISQUE [Mittal et al., 2012]. Consistency with the LR input is measured via peak Signal-to-Noise Ratio (PSNR, denoted as LR-PSNR in Tables 1 and 2). Furthermore, to evaluate the diversity of the super-resolution, we compute the Average Pairwise distance between different samples coming from the same LR input (denoted as APD in Tables 1 and 2), both at the pixel level, using the mean square error (MSE) between samples (considering pixel intensity value between 0 and 1), and at a perceptual level using the LPIPS similarity criteria [Zhang et al., 2018]. For one LR input, the average pairwise distance is computed as the average distance between all the possible pairs of images in a set of 5 super-resolved samples. The reported APD in Tables 1 and 2 corresponds to the mean value of the single image APD over 500 LR inputs in the test set. We measure the distortion of the super-resolved samples with respect to the ground truth HR image in terms of PSNR, structural similarity (SSIM) [Wang et al., 2004] and LPIPS, as it is common in the super-resolution literature. All numbers reported correspond to the metric mean value on a subset of 1000 images from FFHQ256 test set.
@@ 4.2 Results
The quantitative results presented in Table 1 indicate that CVDVAE provides a good trade-off between the different evaluated metrics. Indeed, it obtains the second best results in terms of distortion and visual quality, and the second or third best results in terms of diversity. CVDVAE is also one of the fastest methods, along with HCFlow. HCFlow provides the best results for distortion metrics as it explicitly penalizes bad reconstruction in its training loss. Similar to CVDVAE, its application is fast, as it requires only one network evaluation to produce a super-resolved image. However, HCFlow lacks high-level diversity (as measured by the LPIPS average pairwise distance), compared with the concurrent methods. We postulate that this lack of diversity is due to the relative lack of expressiveness of normalizing flows architecture compared to the convolutional architectures used by diffusion and HVAE models. Our method, along with DPS, produces the best results in terms of visual quality as measured by the BRISQUE metric, illustrating the benefit of using a pretrained unconditional generative model. The computational cost of DPS is nevertheless significantly higher than the ones of CVDVAE and HCFlow, as DPS requires $1000$ steps of network evaluations (and backpropagation through the denoiser) to produce one super-resolved sample. Finally, SR3 performances are inferior to the compared methods. We used the same computational budget (48 hours on 4 GPUs) for training the SR3 models than our CVDVAE and HCFlow. This computational budget is significantly lower than the one reported in the SR3 paper [Saharia et al., 2021] ($\approx 4$ days on 64 TPUv3 chip), and we expect that training the SR3 model for more epochs would improve its performance. Like DPS, SR3 is slower than our method as it requires 2000 network evaluations to produce one super-resolved image; although, unlike DPS, SR3 does not require to backpropagate through the score-network.
@@
Temperature control As for the unconditionnal HVAE models, CVDVAE offers the possibility to control the conditional generation via the temperature of the latent variable distributions [Vahdat and Kautz, 2020, Child, 2021]. The temperature parameter $\tau$ controls the variance of the Gaussian latent distributions. In order to assess the behavior of the model on both low and high temperature regime, we evaluate our method on 2 temperatures ($\tau\in\{0.1,0.8\}$). Quantitative results in Table 2 show that reducing the temperature leads to a solution closer to the ground truth in terms of low-levels distortion metrics (PSNR and the SSIM), while using a higher temperature helps to improve the perceptual similarity (LPIPS) with the ground-truth, as well as the general perceptual quality of the generated HR images and the diversity of the samples. On Figure 5, we display CVDVAE’s samples at different temperatures $\tau$. The sampling temperature correlates with the perceptual smoothness of the super-resolved sample, a higher sampling temperature inducing images with sharper details.
@@ 5 RELATED WORKS
A large number of methods were designed to solve imaging inverse problems such as image super-resolution by using pretrained deep generative models (DGM) as a prior. This includes methods relying on generative adversarial networks (GAN) [Menon et al., 2020, Marinescu et al., 2020, Pan et al., 2021, Daras et al., 2021, Daras et al., 2022, Poirier-Ginter and Lalonde, 2023], variational autoencoders [Mattei and Frellsen, 2018, González et al., 2021, Prost et al., 2023] and denoising diffusion models [Choi et al., 2021, Chung et al., 2022, Kawar et al., 2022, Song et al., 2023]. However, those approaches are computationally expensive as they require an iterative sampling or optimization procedure which require many network evaluation. On the other hand, our approach enables fast inference (one network evaluation), at the cost of reduced flexibility (due to the need of training a task-specific encoder). The idea of training an encoder to map a degraded image in the latent space of a generative network was previously exploited in the context of image inpainting with HVAE [Harvey et al., 2022], and super-resolution with a GAN prior [Chan et al., 2021, Richardson et al., 2021].
@@
Although it is possible to sample from the posterior $p(\bm{x}|\bm{y})$ by using an unconditionnal deep generative models, those methods are restricted to specific dataset for which pretrained models are available. On generic natural images, the state of the art methods rely on conditional generative models, directly trained to model the posterior $p(\bm{x}|\bm{y})$ [Lugmayr et al., 2021, Lugmayr et al., 2022]. Those methods include conditional normalizing flows [Lugmayr et al., 2020, Liang et al., 2021], conditional GAN [Bahat and Michaeli, 2020], conditional VAE [Gatopoulos et al., 2020, Zhou et al., 2021, Chira et al., 2022] and conditional denoising diffusion models [Li et al., 2022, Saharia et al., 2021].
@@ 6 CONCLUSIONS
Our results are in line with many works that illustrates the benefits of using HVAE models for downstream applications [Havtorn et al., 2021, Agarwal et al., 2023, Prost et al., 2023]. One drawback of our approach is its limitation to dataset for which pretrained HVAE models are available, such as human faces or low-resolution ImageNet. However, we postulate that HVAE models have not yet reached their limits, and, by adapting design features from current SOTA deep generative models [Rombach et al., 2022, Kang et al., 2023] (architectural improvement, longer training, larger dataset), HVAE models could significantly improve their performance and expreessiveness, and generalize on much diverse datasets.

### 09 Jan 2024 v4

@@ 1 INTRODUCTION
Image super-resolution is the task of generating a high-resolution (HR) image $\bm{x}$ corresponding to a low-resolution (LR) observation $\bm{y}$. A typical approach for image super-resolution is to train a deep neural network in a supervised fashion to map a LR image to its HR counterpart (see [23] for an extensive review). Despite impressive performances, those regression based methods are fundamentally limited by their lack of diversity. Indeed, there might exist many plausible HR solutions associated with one LR observation, but regression based methods only provide one of those solutions.
@@
Starting from the seminal work of  [26], many approaches proposed to train conditional generative models such as conditional normalizing flow or conditional variational autoencoders in order to model the posterior distribution of the super-resolution problem. We refer to those methods as direct methods, as they only require one network function evaluation (NFE) to generate one sample.
With the recent development of score-based generative models (also known as denoising diffusion models) [16, 41], posterior sampling methods based on conditional denoising diffusion models are now able to produce high-quality samples outperforming previous direct methods [6, 7, 20]. However, denoising diffusion methods are limited by their computationally expensive sampling process, as they require numerous network function evaluations to generate one super-resolved sample. In the following, we classify those methods as iterative methods.
@@
Our approach is based on reusing a pretrained hierarchical variational autoencoder (HVAE) [39, 21]. HVAE models are able to generate high-quality images by relying on an expressive sequential generative model [42, 4, 15]. By associating one latent variable subgroup to each residual block of a generative network, HVAE models are able to learn compact high-level representations of the data, and they can generate new samples efficiently, with only one evaluation of the generative network.
The fast sampling time and the expressivity of HVAE models make them suitable candidates for efficient posterior sampling. In this work, we exploit a pretrained VD-VAE model [4]. In order to repurpose VD-VAE generative model for image super-resolution, we train a low-resolution encoder to encode LR images in the latent space of the VD-VAE model. By combining the LR encoder with the VD-VAE generative model, we can produce a sample with only one (autoencoder) network evaluation. By adopting a stochastic model for the LR encoder, our method can generate diverse samples from the posterior distribution, as illustrated in Figure 1. We show that the LR encoder can be trained with reasonable computational resources by exploiting the VD-VAE original (HR) encoder to generate labels for training the LR-encoder, and by sharing weights between the LR encoder and VD-VAE generative model. We evaluate our method on super-resolution of face images, with upscaling factor ×$\times 4$ and ×$\times 8$ and demonstrate that it reaches sample quality on par with sequential methods, while being significantly faster (>×500$\times$ 500> × 500).
@@ 2 HIERARCHICAL VAE
A VAE also provides an inference model (encoder) $q_{\phi}\left(\bm{z}|\bm{x}\right)$, trained to match the intractable model posterior $p_{\theta}\left(\bm{z}|\bm{x}\right)$ [22]. In order to define expressive models, both the generative model $p_{\theta}\left(\bm{z},\bm{x}\right)$ and the encoder $q_{\phi}\left(\bm{z}|\bm{x}\right)$ are parameterized by neural networks, whose weights are respectively parameterized with $\theta$ and $\phi$.
@@ 3 SUPER-RESOLUTION WITH HVAE
Super-resolution model It has been observed in several works that the low-frequency information of images generated by HVAE model where mostly controlled by the low-resolution latent variable, at the beginning of the hierarchy [42, 4, 14]. Preliminary experiments, detailed in appendix A validate those previous observations. Hence, for a large enough number of latent groups $k$, samples from $p_{\theta}\left(\bm{x}|\bm{x}_{<k}\right)$ share the same low-frequency information. As a consequence, all the samples from $p_{\theta}\left(\bm{x}|\bm{x}_{<k}\right)$ are consistent to a LR image $\bm{y}$ (up to a small error). This motivates us to define the following super-resolution model:
@@
The criterion (13) was introduced by [13], who demonstrated that minimizing (13) is equivalent to maximizing a lower-bound of the super-resolution conditional log-likelihood on the training dataset, and that under additional assumptions on the pretrained HVAE model, one can reach optimal performance by only training the low-resolution encoder $q_{\psi}\left(\bm{z}_{<k}|\bm{y}\right)$. In practice, the KL divergence within the training criterion (13) can be decomposed into a sum of KL divergence on each latent subgroup:
@@ 4.1 Experimental settings
We test our super-resolution method CVDVAE on the FFHQ dataset [19], with images of resolution $256\times 256$. We experiment on 2 upscaling factors: ×$\times 4$ ($64\times 64\rightarrow 256\times 256$) and ×$\times 8$ ($32\times 32\rightarrow 256\times 256$). The low resolution images are initially downscaled by applying an antialiasing kernel followed by a bicubic interpolation.
Compared methods We compare CVDVAE with a conditional normalizing flow (HCFlow) [25], a conditional diffusion model (SR3) [38], and a method that add guidance to a non-conditional diffusion model at inference (DPS) [7]. We retrain HCFlow on FFHQ256 using the official implementation. For DPS, we also reuse the official implementation with the available pretrained model, which was trained on FFHQ. For SR3, since no official implementation is available, we used an open-source (non-official) implementation [17], and we trained a model on FFHQ. When training SR3, we found that a color shift [10] was responsible for important reconstruction errors. To compensate this weakness of the method, we project the super-resolved image on the space of consistent solutions at inference as proposed in [2]. For fair comparison, we retrained both HCFlow and SR3 with the same computational budget as for our low-resolution encoder. For HCFlow and CVDVAE, we set the temperature of the latent variables at $\tau=0.8$ during sampling.
@@
Evaluation metrics The general quality of the super-resolved images is evaluated using the blind Image quality metric BRISQUE [32]. Consistency with the LR input is measured via peak Signal-to-Noise Ratio (PSNR, denoted as LR-PSNR in Tables 1 and 2). Furthermore, to evaluate the diversity of the super-resolution, we compute the Average Pairwise distance between different samples coming from the same LR input (denoted as APD in Tables 1 and 2), both at the pixel level, using the mean square error (MSE) between samples (considering pixel intensity value between 0 and 1), and at a perceptual level using the LPIPS similarity criteria [44]. For one LR input, the average pairwise distance is computed as the average distance between all the possible pairs of images in a set of 5 super-resolved samples. The reported APD in Tables 1 and 2 corresponds to the mean value of the single image APD over 500 LR inputs in the test set. We measure the distortion of the super-resolved samples with respect to the ground truth HR image in terms of PSNR, structural similarity (SSIM) [43] and LPIPS, as it is common in the super-resolution literature. All numbers reported correspond to the metric mean value on a subset of 1000 images from FFHQ256 test set.
@@ 4.2 Results
The quantitative results presented in Table 1 indicate that CVDVAE provides a good trade-off between the different evaluated metrics. Indeed, it obtains the second best results in terms of distortion and visual quality, and the second or third best results in terms of diversity. CVDVAE is also one of the fastest methods, along with HCFlow. HCFlow provides the best results for distortion metrics as it explicitly penalizes bad reconstruction in its training loss. Similar to CVDVAE, its application is fast, as it requires only one network evaluation to produce a super-resolved image. However, HCFlow lacks high-level diversity (as measured by the LPIPS average pairwise distance), compared with the concurrent methods. We postulate that this lack of diversity is due to the relative lack of expressiveness of normalizing flows architecture compared to the convolutional architectures used by diffusion and HVAE models. Our method, along with DPS, produces the best results in terms of visual quality as measured by the BRISQUE metric, illustrating the benefit of using a pretrained unconditional generative model. The computational cost of DPS is nevertheless significantly higher than the ones of CVDVAE and HCFlow, as DPS requires $1000$ steps of network evaluations (and backpropagation through the denoiser) to produce one super-resolved sample. Finally, SR3 performances are inferior to the compared methods. We used the same computational budget (48 hours on 4 GPUs) for training the SR3 models than our CVDVAE and HCFlow. This computational budget is significantly lower than the one reported in the SR3 paper [38] ($\approx 4$ days on 64 TPUv3 chip), and we expect that training the SR3 model for more epochs would improve its performance. Like DPS, SR3 is slower than our method as it requires 2000 network evaluations to produce one super-resolved image; although, unlike DPS, SR3 does not require to backpropagate through the score-network.
@@
Temperature control As for the unconditionnal HVAE models, CVDVAE offers the possibility to control the conditional generation via the temperature of the latent variable distributions [42, 4]. The temperature parameter $\tau$ controls the variance of the Gaussian latent distributions. In order to assess the behavior of the model on both low and high temperature regime, we evaluate our method on 2 temperatures ($\tau\in\{0.1,0.8\}$). Quantitative results in Table 2 show that reducing the temperature leads to a solution closer to the ground truth in terms of low-levels distortion metrics (PSNR and the SSIM), while using a higher temperature helps to improve the perceptual similarity (LPIPS) with the ground-truth, as well as the general perceptual quality of the generated HR images and the diversity of the samples. On Figure 5, we display CVDVAE’s samples at different temperatures $\tau$. The sampling temperature correlates with the perceptual smoothness of the super-resolved sample, a higher sampling temperature inducing images with sharper details.
@@ 5 RELATED WORKS
A large number of methods were designed to solve imaging inverse problems such as image super-resolution by using pretrained deep generative models (DGM) as a prior. This includes methods relying on generative adversarial networks (GAN) [31, 29, 33, 9, 8, 34], variational autoencoders [30, 12, 35] and denoising diffusion models [6, 7, 20, 40]. However, those approaches are computationally expensive as they require an iterative sampling or optimization procedure which require many network evaluation. On the other hand, our approach enables fast inference (one network evaluation), at the cost of reduced flexibility (due to the need of training a task-specific encoder). The idea of training an encoder to map a degraded image in the latent space of a generative network was previously exploited in the context of image inpainting with HVAE [13], and super-resolution with a GAN prior [3, 36].
@@
Although it is possible to sample from the posterior $p(\bm{x}|\bm{y})$ by using an unconditionnal deep generative models, those methods are restricted to specific dataset for which pretrained models are available. On generic natural images, the state of the art methods rely on conditional generative models, directly trained to model the posterior $p(\bm{x}|\bm{y})$ [27, 28]. Those methods include conditional normalizing flows [26, 25], conditional GAN [2], conditional VAE [11, 45, 5] and conditional denoising diffusion models [24, 38].
@@ 6 CONCLUSIONS
Our results are in line with many works that illustrates the benefits of using HVAE models for downstream applications [14, 1, 35]. One drawback of our approach is its limitation to dataset for which pretrained HVAE models are available, such as human faces or low-resolution ImageNet. However, we postulate that HVAE models have not yet reached their limits, and, by adapting design features from current SOTA deep generative models [37, 18] (architectural improvement, longer training, larger dataset), HVAE models could significantly improve their performance and expreessiveness, and generalize on much diverse datasets.